# TacMan-Turbo: Proactive Tactile Control for Robust and Efficient Articulated Object Manipulation

### v2

@@ 3.1 Notation
  • The unit vector $\hat{\boldsymbol{\omega}}\in\mathbb{R}^{3}$ specifies the axis of rotation, defining the direction of rotational motion for articulated objects.
3.2 Problem Formulation
We now formulate the mathematical framework for effective and efficient tactile-informed manipulation of articulated objects. Consider an articulated object with a movable part and its rigidly attached “handle,” as illustrated in Fig. 2. The underlying articulation mechanism constrains the handle’s motion to a 1-DoF trajectory, represented as path $J$ from initial point $J_{s}$ to target point $J_{e}$. In our problem setting, only the initial position $J_{s}$ is initially known, reflecting real-world scenarios where complete kinematic models are unavailable.
@@ 3.2 Problem Formulation
The physical interaction between the gripper and handle is characterized by their contact state, denoted by $\mathcal{C}_{i}$. This state is represented geometrically as a set containing three-dimensional position vectors of activated contact points in the gripper’s coordinate frame, denoted as $p_{x}^{g_{i}},p_{y}^{g_{i}},p_{z}^{g_{i}}]^{\mathsf{T}}$. Contact points become activated when their normal deformation (illustrated along the $x$-axis in Fig. 2) exceeds threshold $\epsilon$:
$\mathcal{C}_{i}=\left\{\begin{bmatrix}p_{x}^{g_{i}}\\ p_{y}^{g_{i}}\\ p_{z}^{g_{i}}\end{bmatrix}\;\middle|\;\lvert p_{x}^{g_{i}}\rvert\geq\epsilon\right\}.$ (2)
@@
Manipulation begins with a constant base velocity $\boldsymbol{u}_{0}^{g_{i}}$ in the gripper’s coordinate system, aligned with a known rough direction of the manipulation task:
$\boldsymbol{u}_{0}^{g_{i}}=\begin{bmatrix}\boldsymbol{v}_{0}^{g_{i}}\\ \boldsymbol{\omega}_{0}^{g_{i}}\end{bmatrix}$. (4)
The acquisition of this base velocity and its impact on manipulation efficiency are examined in detail in Section 6.
@@
Given that the contact state depends solely on the poses of the gripper and handle (as established in Eq. 3), maintaining stable contact requires aligning the gripper’s motion with the handle’s trajectory. Our solution addresses this fundamental challenge in two complementary steps: First, in Section 3.3, we develop a method to predict the handle’s next-step pose $\boldsymbol{T}_{h_{i+1}}$ based on sequential contact patterns observed during interaction. Second, in Section 3.4, we compute the control input $\boldsymbol{u}_{i}^{g_{i}}$ that optimally aligns the gripper’s motion with this predicted pose while ensuring compliance with the velocity limit constraints.
3.3 Handle Pose Estimation and Prediction
@@
where $\boldsymbol{T}^{g_{i}}_{h_{i}}\in SE(3)$ represents their relative homogeneous transformation. Initially at $i=1$, $\boldsymbol{T}^{g_{1}}_{h_{1}}=\boldsymbol{I}_{4\times 4}$ following our initial pose definition (Eq. 1).
@@
Building upon these estimated poses, we now focus on predicting future handle positions. Under the assumptions of smooth object trajectories and sufficiently high sampling frequencies, consecutive handle poses exhibit only slight variations primarily driven by the object’s instantaneous velocity. This makes the constant velocity model—a widely utilized approach across various scenarios, including dynamic ones [11, 43, 48]—particularly effective for such predictions. Based on this model, the next handle pose ${\tilde{\boldsymbol{T}}}_{h_{i+1}}$ is predicted as:
@@
As evident from Eq. 15, the proposed constant velocity prediction requires at least two previous observations ($i\geq 2$). For the initial step ($i=1$), when past data is unavailable, the system relies solely on the base velocity $\boldsymbol{u}_{0}^{g_{1}}$ defined in Section 3.2. This integration of pose estimation and prediction provides the foundation for computing appropriate gripper velocities, which we address in the following section.
3.4 Offset Velocity Computation
@@
This desired alignment requires the gripper to execute a specific transformation over a sampling interval $\Delta t\in\mathbb{R^{+}}$. We compute this required transformation relative to the gripper’s current coordinate frame:
@@
From this transformation, we derive the required velocity components. The linear velocity component $\boldsymbol{v}_{i}^{g_{i}}$ of the offset velocity $\boldsymbol{u}_{i}^{g_{i}}$ is computed by accounting for both the position difference and the base velocity:
$\boldsymbol{v}_{i}^{g_{i}}=\frac{\breve{\boldsymbol{p}}_{i}^{g_{i}}}{\Delta t}-\boldsymbol{v}_{0}^{g_{i}}.$ (18)
For the angular velocity component $\boldsymbol{\omega}_{i}^{g_{i}}$, we employ Rodrigues’ rotation formula [37]:
$\boldsymbol{\omega}_{i}^{g_{i}}=\hat{\boldsymbol{\omega}}_{i}^{g_{i}}\frac{\theta_{i}}{\Delta t}-\boldsymbol{\omega}_{0}^{g_{i}},$ (19)
where $\theta_{i}\in[0,\pi]$ represents the rotation angle:
$\theta_{i}=\cos^{-1}\left(\frac{1}{2}(\text{tr}(\breve{\boldsymbol{R}}_{i}^{g_{i}})-1)\right),$ (20)
and $\hat{\boldsymbol{\omega}}_{i}^{g_{i}}\in\mathbb{R}^{3}$ denotes the unit vector along the rotation axis, whose computation depends on the rotation angle $\theta_{i}$. For brevity, the details of this computation are omitted here; interested readers are referred to the comprehensive treatment in [37].
With both linear and angular components determined, we can express the ideal offset velocity:
$\boldsymbol{u}_{i}^{g_{i}}=\begin{bmatrix}\boldsymbol{v}_{i}^{g_{i}}\\ \boldsymbol{\omega}_{i}^{g_{i}}\end{bmatrix}$. (21)
While the above formulation provides the theoretically optimal velocity, practical implementation must account for the limitations imposed by robot hardware—specifically the maximum gripper velocities determined by the robot’s joint motor limits, as defined in Eq. 7. The relationship between joint velocities $\dot{\boldsymbol{q}}_{i}\in\mathbb{R}^{n}$ and the gripper velocity $\boldsymbol{u}_{i}^{g_{i}}$ is given by:
@@
$\alpha=\min\left(\min\left(\left\{\frac{\dot{q}_{M_{j}}}{\max(\lvert\dot{q}_{i_{j}}\rvert,\epsilon)}\beta\;\middle|\;j=1,\ldots,n\right\}\right),1\right),$ (25)
where $\beta$ is a safety factor and $\epsilon$ represents a small positive constant introduced to prevent division by zero. This scaling yields the final executable offset velocity command:
$\boldsymbol{u}_{i}^{g_{i}}=\begin{bmatrix}\alpha\boldsymbol{v}_{i}^{g_{i}}+(\alpha-1)\boldsymbol{v}_{0}^{g_{i}}\\ \alpha\boldsymbol{\omega}_{i}^{g_{i}}+(\alpha-1)\boldsymbol{\omega}_{0}^{g_{i}}\end{bmatrix}$ (26)
@@
Figure 3: Simulation test environment and object categories. Our comprehensive evaluation framework encompasses three progressively complex articulated object categories: (a) 50 prismatic-joint objects featuring linear motion paths, representing the most kinematically straightforward mechanisms; (b) 50 revolute-joint objects with circular motion paths, introducing rotational complexity; and (c) 100 complex articulated objects with trajectories derived from Bézier curves of orders 2–5, simulating sophisticated real-world mechanisms with variable curvature. This systematically designed test suite enables rigorous performance assessment across a spectrum of kinematic complexity. Each case is standardized using initial robot configurations, ensuring fair comparison across manipulation strategies.
4 Simulation Studies
To thoroughly evaluate TacMan-Turbo, we conducted extensive simulation studies focusing on both manipulation effectiveness and efficiency. This section presents our experimental approach and findings in four parts: we first detail our simulation setup (Section 4.1), establish comparative baselines (Section 4.2), define quantitative evaluation metrics (Section 4.3), and analyze the performance results (Section 4.4).
4.1 Simulation Setup
@@
We implemented our experiments in NVIDIA Isaac Sim, operating at $60~\mathrm{Hz}$ to ensure sufficient temporal resolution for capturing dynamic contact interactions. Critical to our study is the contact simulation method developed by Zhao et al[73], which enables accurate modeling of tactile feedback by monitoring displacement patterns within designated contact regions. This approach captures the nuanced contact dynamics that emerge during the manipulation of articulated objects.
@@
  • 50 objects with prismatic joints from PartNet-Mobility dataset [60] (Fig. 3(a)), featuring linear motion paths with varying physical characteristics.
  • 50 objects with revolute joints also from PartNet-Mobility (Fig. 3(b)), representing rotational articulations that require continuous adjustment of manipulation direction.
  • 100 custom-generated objects with Bézier curve trajectories designed to assess performance on complex, non-linear paths. This category includes 25 curves each of orders 2, 3, 4, and 5, sampled on a 2D plane and constrained to avoid self-intersections. Each curve defines a manipulation path on a 3D playboard, with a toy train positioned at the starting point serving as the handle (Fig. 3(c)).
Such a design from simple to complex kinematics provides systematic evaluation across a spectrum of trajectory predictability and curvature characteristics. The generation process for the Bézier-based objects followed protocols in Zhao et al[73], producing diverse trajectory complexities that challenge predictive capabilities in distinct ways.
@@
4.2 Baseline Selection
For comparative analysis, we adopt Tac-Man [73] as our primary baseline, a selection driven by its demonstrated superiority in manipulation effectiveness, outperforming traditional pre-planned methods, impedance control strategies, and learning-based approaches across diverse manipulation scenarios. Since successful task completion is a prerequisite for meaningful efficiency evaluation, we focus our comparative analysis exclusively on methods that reliably satisfy this fundamental criterion.
@@ 4.2 Baseline Selection
The key architectural distinction between Tac-Man and our proposed TacMan-Turbo framework lies in their fundamental approach to manipulation control. Tac-Man implements a two-stage reactive paradigm: an execution stage that performs the primary manipulation, followed by a separate recovery stage that corrects contact deviations after they occur. In contrast, TacMan-Turbo integrates these functions into a unified proactive process that anticipates and preemptively addresses potential deviations, enabling more fluid and continuous manipulation.
@@
To ensure a fair and rigorous comparison, both methods operate under identical initialization conditions as specified in Section 4.1, including consistent robot configuration, initial grasp state, and interaction direction. This standardized experimental protocol effectively isolates the impact of their differing control strategies, ensuring that performance differences directly reflect the fundamental distinction between reactive and proactive approaches to manipulation.
@@
Figure 4: Qualitative results of simulation studies. This visualization demonstrates TacMan-Turbo’s performance across three representative cases from our object categories: prismatic (left), revolute (middle), and complex articulation (right). The top row displays trajectory tracking performance with actual handle positions (blue), predicted handle positions (orange), gripper positions (green), and contact points (red), while the bottom row shows corresponding tactile feedback patterns. For objects with consistent kinematic patterns (prismatic and revolute), TacMan-Turbo’s prediction mechanism generates remarkably accurate estimates with minimal residual error, enabling precise trajectory alignment. When confronting complex articulations with sudden directional changes (right column), prediction accuracy temporarily decreases at inflection points, but the system rapidly converges back to the optimal path through iterative correction. The magnified insets highlight these prediction characteristics, demonstrating how TacMan-Turbo maintains effective manipulation even when faced with kinematically challenging scenarios. Contact deviation vectors (orange arrows, magnified 5× for visibility) further illustrate how the system continuously adjusts to maintain optimal contact.
4.3 Evaluation Metrics
@@
As our primary measure of manipulation effectiveness, success rate employs clear binary criteria tailored to each object type. For articulated objects with standard joints, success requires either rotating a revolute joint beyond $60.0^{\circ}$ or extending a prismatic joint beyond $250.0~\mathrm{mm}$. For the Bézier curve trajectories, success is achieved when the gripper successfully guides the toy train from one end of the track to the other. Manipulations failing to meet these thresholds are classified as unsuccessful, regardless of progress made.
@@
Action efficiency evaluates how directly each motion contributes to task progress, providing insights into the manipulation strategy’s precision and economy of movement. We implement joint-specific calculations: for revolute joints, we compute normalized angular displacement relative to a $60^{\circ}$ threshold; for prismatic joints, we measure normalized linear displacement relative to a $250~\mathrm{mm}$ threshold.
@@
4.4 Simulation Results
We begin by presenting qualitative results of TacMan-Turbo’s performance in manipulating articulated objects. Using representative examples from each category described in Section 4.1, Fig. 4 illustrates how the predicted handle poses effectively guide the gripper during manipulation tasks. Our observations reveal that when object motion exhibits smooth and consistent trends, the prediction model captures this behavior with high accuracy. However, as shown in the rightmost column of Fig. 4, abrupt changes in object motion can occasionally introduce prediction errors, stemming from the non-causal nature of our prediction model. Nevertheless, the system demonstrates robust performance through its iterative prediction process, which continually incorporates new observations to correct these errors. These results highlight TacMan-Turbo’s practical effectiveness in handling complex manipulation scenarios commonly encountered in real-world applications. Additional qualitative results are available in the Supplementary Video S1.
Our quantitative analysis evaluates TacMan-Turbo against Tac-Man using the four metrics defined in Section 4.3.
Table 1: Success rate (%) comparing Tac-Man and TacMan-Turbo across articulated object categories. Both methods achieve perfect manipulation success across all tested joint types, from simple mechanical joints to complex Bézier curves (BC-n) of increasing order.
Joint type Pri. Rev. BC-2 BC-3 BC-4 BC-5
Tac-Man 100 100 100 100 100 100
TacMan-Turbo 100 100 100 100 100 100

### v3

@@ III-A Notation
III-B Problem Formulation
We now formulate the mathematical framework for effective and efficient tactile-informed manipulation of articulated objects. Consider an articulated object with a movable part and its rigidly attached “handle,” as illustrated in Fig.˜1. The underlying articulation mechanism constrains the handle’s motion to a 1- degree of freedom (DoF) trajectory, represented as path $J$ from initial point $J_{s}$ to target point $J_{e}$. In our problem setting, only the initial position $J_{s}$ is initially known, reflecting real-world scenarios where complete kinematic models are unavailable.
@@ III-B Problem Formulation
The physical interaction between the gripper and handle is characterized by their contact state, denoted by $\mathcal{C}_{i}$. This state is represented geometrically as a set containing three-dimensional position vectors of activated contact points in the gripper’s coordinate frame, denoted as $p_{x}^{g_{i}},p_{y}^{g_{i}},p_{z}^{g_{i}}]^{\mathsf{T}}$. Contact points become activated when their normal deformation (illustrated along the $x$-axis in Fig.˜1) exceeds threshold $\epsilon$:
$\mathcal{C}_{i}=\left\{\begin{bmatrix}p_{x}^{g_{i}}\\ p_{y}^{g_{i}}\\ p_{z}^{g_{i}}\end{bmatrix}\;\middle|\;\lvert p_{x}^{g_{i}}\rvert\geq\epsilon\right\},$ (2)
where the threshold $\epsilon$ is adaptively selected in experiments, with implementation details provided in Sec.˜V-A.
@@
Manipulation begins with a constant base velocity $\boldsymbol{u}_{0}^{g_{i}}$ in the gripper’s coordinate system, aligned with a known rough direction of the manipulation task. The acquisition of this base velocity and its impact on manipulation efficiency are examined in detail in Sec.˜VI.
@@
Given that the contact state depends solely on the poses of the gripper and handle (as established in Eq.˜3), maintaining stable contact requires aligning the gripper’s motion with the handle’s trajectory. Our solution addresses this fundamental challenge in two complementary steps: First, in Sec.˜III-C, we develop a method to predict the handle’s next-step pose $\boldsymbol{T}_{h_{i+1}}$ based on sequential contact patterns observed during interaction. Second, in Sec.˜III-D, we compute the control input $\boldsymbol{u}_{i}^{g_{i}}$ that optimally aligns the gripper’s motion with this predicted pose while ensuring compliance with the velocity limit constraints.
III-C Handle Pose Estimation and Prediction
@@
where $\boldsymbol{T}^{g_{i}}_{h_{i}}\in SE(3)$ represents their relative homogeneous transformation. Initially at $i=1$, $\boldsymbol{T}^{g_{1}}_{h_{1}}=\boldsymbol{I}_{4\times 4}$ following our initial pose definition (Eq.˜1).
@@
Building upon these estimated poses, we now focus on predicting future handle positions. Under the assumptions of smooth object trajectories and sufficiently high sampling frequencies, consecutive handle poses exhibit only slight variations, primarily driven by the object’s instantaneous velocity. This makes the constant velocity model—a widely utilized approach across various scenarios, including dynamic ones [86, 87, 88]—particularly effective for such predictions. Based on this model, the next handle pose ${\tilde{\boldsymbol{T}}}_{h_{i+1}}$ is predicted as:
@@
As evident from Eq.˜14, the proposed constant velocity prediction requires at least two previous observations ($i\geq 2$). For the initial step ($i=1$), when past data is unavailable, the system relies solely on the base velocity $\boldsymbol{u}_{0}^{g_{1}}$ defined in Sec.˜III-B. This integration of pose estimation and prediction provides the foundation for computing appropriate gripper velocities, which we address in the following section.
III-D Offset Velocity Computation
@@
This desired alignment requires the gripper to execute a specific transformation over a sampling interval $\Delta t\in\mathbb{R}_{>0}$. We compute this required transformation relative to the gripper’s current coordinate frame:
@@
From this transformation, we derive the required velocity components:
$\boldsymbol{u}_{i}^{g_{i}}=\frac{\log(\breve{\boldsymbol{T}}_{i}^{g_{i}})^{\vee}}{\Delta t}-\boldsymbol{u}_{0}^{g_{i}},$ (17)
where $\log(\cdot)$ denotes the logarithmic map and the operator $\cdot^{\vee}$ transforms the skew-symmetric matrix into vector form.
While the above formulation provides the theoretically optimal velocity, practical implementation must account for the limitations imposed by robot hardware—specifically the maximum gripper velocities determined by the robot’s joint motor limits, as defined in Eq.˜6. The relationship between joint velocities $\dot{\boldsymbol{q}}_{i}\in\mathbb{R}^{n}$ and the gripper velocity $\boldsymbol{u}_{i}^{g_{i}}$ is given by:
@@
$\alpha=\min\left(\min\left(\left\{\frac{\dot{q}_{M_{j}}}{\max(\lvert\dot{q}_{i_{j}}\rvert,\delta)}\beta\;\middle|\;j=1,\ldots,n\right\}\right),1\right),$ (21)
where $\beta$ is a safety factor and $\delta$ represents a small positive constant introduced to prevent division by zero. In this study, we select $\beta=2$ and $\delta=1\times 10^{-4}$ across all the experiments. This scaling yields the final executable offset velocity command:
$\boldsymbol{u}_{i}^{g_{i}}=\alpha\boldsymbol{u}_{i}^{g_{i}}+(\alpha-1)\boldsymbol{u}_{0}^{g_{i}}$ (22)
@@
Figure 2: Simulation test environment and object categories. Our evaluation encompasses three progressively complex articulated object categories: (a) 50 prismatic-joint objects featuring linear motion paths, representing the most kinematically straightforward mechanisms; (b) 50 revolute-joint objects with circular motion paths, introducing rotational complexity; and (c) 100 complex articulated objects with trajectories derived from Bézier curves of orders 2–5, simulating sophisticated real-world mechanisms with variable curvature. This systematically designed test suite enables rigorous performance assessment across a spectrum of kinematic complexity. Each case is standardized using initial robot configurations, ensuring fair comparison across manipulation strategies.
IV Simulation Studies
To thoroughly evaluate TacMan-Turbo, we conducted extensive simulation studies focusing on both manipulation effectiveness and efficiency. This section presents our experimental approach and findings in four parts: we first detail our simulation setup (Sec.˜IV-A), establish comparative baselines (Sec.˜IV-B), define quantitative evaluation metrics (Sec.˜IV-C), and analyze the performance results (Sec.˜IV-D).
IV-A Simulation Setup
@@
We implemented our experiments in NVIDIA Isaac Sim, operating at $60\text{\,}\mathrm{Hz}$ to ensure sufficient temporal resolution for capturing dynamic contact interactions. Critical to our study is the contact simulation method developed by Zhao et al. [1], which enables accurate modeling of tactile feedback by monitoring displacement patterns within designated contact regions. This approach captures the nuanced contact dynamics that emerge during the manipulation of articulated objects.
@@
  • 50 objects with prismatic joints from PartNet-Mobility dataset [40] (Fig.˜2(a)), featuring linear motion paths with varying physical characteristics.
  • 50 objects with revolute joints also from PartNet-Mobility (Fig.˜2(b)), representing rotational articulations that require continuous adjustment of manipulation direction.
  • 100 custom-generated objects with Bézier curve trajectories designed to assess performance on complex, non-linear paths. This category includes 25 curves each of orders 2, 3, 4, and 5, sampled on a 2D plane and constrained to avoid self-intersections. Each curve defines a manipulation path on a 3D playboard, with a toy train positioned at the starting point serving as the handle (Fig.˜2(c)).
Such a design from simple to complex kinematics provides systematic evaluation across a spectrum of trajectory predictability and curvature characteristics. The generation process for the Bézier-based objects followed protocols in Zhao et al. [1], producing diverse trajectory complexities that challenge predictive capabilities in distinct ways.
@@
IV-B Baseline Selection
@@
For comparative analysis, we adopt Tac-Man [1] as our primary baseline, a selection driven by its demonstrated superiority in manipulation effectiveness. The key architectural distinction between Tac-Man and our proposed TacMan-Turbo framework lies in their fundamental approach to manipulation control. Tac-Man implements a two-stage reactive paradigm: an execution stage that performs the primary manipulation, followed by a separate recovery stage that corrects contact deviations after they occur. In contrast, TacMan-Turbo integrates these functions into a unified proactive process that anticipates and preemptively addresses potential deviations, enabling more fluid and continuous manipulation.
Beyond the primary baseline, we evaluate TacMan-Turbo against three additional baselines spanning traditional pre-planned methods, impedance control strategies, and learning-based approaches to contextualize our contribution within the broader literature. The pre-planned method leverages known kinematics to generate trajectories. To model parameter estimation uncertainty in real-world deployment, we perturb the ground-truth kinematic parameters with Gaussian noise (variance equals $0.01$). The impedance control baseline operates under the same prior-free setting as Tac-Man but excludes tactile feedback, instead relying on impedance adaptation to accommodate deviations. Finally, the learning-based baseline employs FlowBot3D [19], which is trained to use visual feedback for manipulation guidance.
@@
To ensure a fair and rigorous comparison, both methods operate under identical initialization conditions as specified in Sec.˜IV-A, including consistent robot configuration, initial grasp state, and interaction direction. This standardized experimental protocol effectively isolates the impact of their differing control strategies, ensuring that performance differences directly reflect the fundamental distinction.
@@
Figure 3: Qualitative results of simulation studies. This visualization demonstrates TacMan-Turbo’s performance across three representative cases from our object categories: prismatic (left), revolute (middle), and complex articulation (right). The top row displays trajectory tracking performance with actual handle positions (blue), predicted handle positions (orange), gripper positions (green), and contact points (red), while the bottom row shows corresponding tactile feedback patterns. For objects with consistent kinematic patterns (prismatic and revolute), TacMan-Turbo’s prediction mechanism generates remarkably accurate estimates with minimal residual error, enabling precise trajectory alignment. When confronting complex articulations with sudden directional changes (right column), prediction accuracy temporarily decreases at inflection points, but the system rapidly converges back to the optimal path through iterative correction. The magnified insets highlight these prediction characteristics, demonstrating how TacMan-Turbo maintains effective manipulation even when faced with kinematically challenging scenarios. Contact deviation vectors (orange arrows, magnified 5× for visibility) further illustrate how the system continuously adjusts to maintain optimal contact. More qualitative results are available in the Supplementary Video S1.
(a) Distribution of mean errors.
(b) Distribution of error standard deviations.
Figure 4: Prediction errors of various predictive models. The prediction error distributions for three kinematic models—constant velocity, constant acceleration, and constant jerk—are presented and compared. All methods demonstrate comparable performance overall. However, higher-order models exhibit superior accuracy under low-noise conditions, as they better capture the underlying motion dynamics. Conversely, these same models show degraded performance in high-noise environments due to the cumulative effects of numerical differentiation, which amplifies measurement noise with each successive derivative computation. Box plots display time distributions with medians (center lines), interquartile ranges (boxes), and distribution extent (whiskers at $1.5\timesIQR$). BC-n represents Bézier curves of order n.
TABLE I: Success rate (%) comparing TacMan-Turbo and baseline methods across articulated object categories. Both methods achieve perfect manipulation success across all tested joint types, from simple mechanical joints to complex Bézier curves (BC-n) of increasing order. Results are reported as mean (standard deviation).
Method Pri. Rev. BC-2 BC-3 BC-4 BC-5
Planing $96$ p m 16 $93$ p m 21 $71$ p m 29 $39$ p m 37 $53$ p m 40 $43$ p m 36
Imped. $81$ p m 33 $48$ p m 28 $32$ p m 28 $32$ p m 22 2020 p m 23 $43$ p m 27
FlowBot $38$ p m 35 $42$ p m 40 99 p m 6 88 p m 5 88 p m 7 88 p m 9
Tac-Man $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0
Ours $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0 $100$ p m 0
IV-C Evaluation Metrics
@@
As our primary measure of manipulation effectiveness, success rate employs a progress-based criterion tailored to each object type. For articulated objects with standard joints, we measure the percentage of the target displacement achieved: revolute joints are evaluated based on progress toward a $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ rotation, while prismatic joints are assessed based on progress toward a $250\text{\,}\mathrm{mm}$ extension. For the Bézier curve trajectories, success is quantified as the percentage of track distance covered by the toy train from start to end.
@@
Action efficiency evaluates how directly each motion contributes to task progress, providing insights into the manipulation strategy’s precision and economy of movement. We implement joint-specific calculations: for revolute joints, we compute normalized angular displacement relative to a $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ threshold; for prismatic joints, we measure normalized linear displacement relative to a $250\text{\,}\mathrm{mm}$ threshold.
@@
IV-D Simulation Results
We begin by presenting qualitative results of TacMan-Turbo’s performance in manipulating articulated objects. Using representative examples from each category described in Sec.˜IV-A, Fig.˜3 illustrates how the predicted handle poses effectively guide the gripper during manipulation tasks. Our observations reveal that when object motion exhibits smooth and consistent trends, the prediction model captures this behavior with high accuracy. However, as shown in the rightmost column of Fig.˜3, abrupt changes in object motion can occasionally introduce prediction errors, stemming from the non-causal nature of our prediction model. Nevertheless, the system demonstrates robust performance through its iterative prediction process, which continually incorporates new observations to correct these errors. These results highlight TacMan-Turbo’s practical effectiveness in handling complex manipulation scenarios commonly encountered in real-world applications. Additional qualitative results are available in the Supplementary Video S1.
To quantitatively assess the robustness and sensitivity of the predictive model, we evaluate the prediction residuals under varying noise levels and prediction models. While Sec.˜III establishes the constant velocity model as one approach, this section extends the analysis to include constant acceleration and constant jerk models.
[…]
The results are shown in Fig.˜4. All models achieve comparable performance overall. As the complexity of object kinematics increases, prediction errors increase correspondingly due to more varied motion patterns. Under low-noise conditions, higher-order models achieve better performance by capturing more complex underlying dynamics. However, they are more sensitive to noise due to the multiple differential computations involved. This tendency is clearly evident for basic prismatic and revolute joint types, but becomes less distinct in more complex kinematic scenarios. These results demonstrate that the constant-velocity model can be a robust choice, aligning with the intuition underlying previous works [86, 87, 88]. While all predictive models inevitably produce prediction residuals, TacMan-Turbo’s high-frequency operation enables rapid detection of deviations and adaptive adjustment. Rather than relying on perfect predictions, the system uses them as initial estimates within a closed-loop control framework that continuously corrects for modeling errors.
Our quantitative analysis evaluates TacMan-Turbo using the four metrics defined in Sec.˜IV-C.